OpenStudy (anonymous):

a recipe that makes 6 servings calls for 1 1/5 cups of juice. how many cups of juice are there per serving?

OpenStudy (anonymous):

7 1/5 cups

OpenStudy (anonymous):



OpenStudy (anonymous):

well to start, lets makethe numbers easier to work with. I'm going to change 1 1/5 cups to 1.2. Now, there are 6 servings in this recipe, so we are just going to drive the amount of juice by six--Does that make sense? So 1.2/6=.2 or 1/5 cups per serving.

OpenStudy (anonymous):

So B :)

OpenStudy (anonymous):

*divide

OpenStudy (anonymous):

thanks I have more can u help me please:)

OpenStudy (anonymous):

Sure thing, what else you got?

OpenStudy (anonymous):

A recipe that makes 6 servings calls for 1/3 cup sugar. Arlene modifies the recipe so that it can serve 9

OpenStudy (anonymous):

oh that makes sense!!!!

OpenStudy (anonymous):

alright, well for this one we need to set up a proportion. The proportion set-up is in the attachment below. You're going to cross multiply, that is, multiply 9 by 1/3 an then divide that value by 6. So, 9 times 1/3 is 3, and 3 divided by 6 is 1/2. So in order to make a recipe that serves 9, she'll need 1/2 a cup of sugar.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

A recipe for 1 lb of trail mix calls for 3.2 oz of cashews. How many ounces of cashews are needed to make 4 lb of trail mix? A. 0.8 oz B. 4.2 oz C. 7.2 oz D. 12.8 oz

OpenStudy (anonymous):

@Tedimariah

OpenStudy (anonymous):

I can try and help if you open a new question. :)

OpenStudy (anonymous):

Sorry, I'm here!

OpenStudy (anonymous):

Alright, we are cross multiplying here again. So every pound calls for 3.2 oz of cashews. So if we set that up in a proportion, x (being the value we want to find) over 4in equivalent to 3.2 over 1 (See attachment). If we multiply 3.2 times 4 and divide that by one we get 12.8 oz. Which is your answer. Or you can look at it logically, and assume that because there are 3.2 oz to a pound, and you want to make four pounds, you just multiply that by 4 getting 12.8 as well.

OpenStudy (anonymous):

thanks Last month, Edwin had $2325.35 that he divided into two bank accounts. He deposited 2/5 into a savings account. He deposited the rest into a checking account. How much money did Edwin deposit into his checking account?

OpenStudy (anonymous):

We need to find what two fifths of 2325.35 is. So this is fairly straight-forward, we are going to divide it by 5 (hence 5ths) and use that value to find the amount deposited into the checking account. So...\[2325.35/5=465.07\]So we know 1/5th of 2325.35 is 465.07. We know that Edwin deposited 2/5ths of that into a savings account, so whatever is left over is what went to the checkings account. So, 2 times 465.07 will get us our savings value:\[465.07*2=930.14\] This is the amount that went into the savings, so now we subtract that from the original amount to find the amount deposited into the checking.\[2325.35-930.14=1395.21\] So he deposited $1395.21 into the checking account.